# Co-AttenDWG: Co-Attentive Dimension-Wise Gating and Expert Fusion for Multi-Modal Offensive Content Detection

### v1

@@ Abstract
Multi-modal learning has become a critical research area because integrating text and image data can significantly improve performance in tasks such as classification, retrieval, and scene understanding. However, despite progress with pre-trained models, current approaches are limited by inadequate cross-modal interactions and static fusion strategies that do not fully exploit the complementary nature of different modalities. To address these shortcomings, we introduce a novel multi-modal Co-AttenDWG architecture that leverages dual-path encoding, co-attention with dimension-wise gating, and advanced expert fusion. Our approach begins by projecting text and image features into a common embedding space, where a dedicated co-attention mechanism enables simultaneous, fine-grained interactions between modalities. This mechanism is further enhanced by a dimension-wise gating network that adaptively regulates the feature contributions at the channel level, ensuring that only the most relevant information is emphasized. In parallel, dual-path encoders refine the representations by processing cross-modal information separately before an additional cross-attention layer further aligns modalities. The refined features are then aggregated via an expert fusion module that combines learned gating and self-attention to produce a robust, unified representation. We validate our approach on the MIMIC and SemEval Memotion 1.0, where experimental results demonstrate significant improvements in cross-modal alignment and state-of-the-art performance, underscoring the potential of our model for a wide range of multi-modal applications.
@@ I Introduction
Multi-modal learning has emerged as a transformative paradigm in artificial intelligence, driven by the need to integrate diverse data sources such as text, images, audio, and video to provide a holistic understanding of complex real-world scenarios [1, 2, 3]. This integration is particularly important in tasks such as classification, sentiment analysis, and information retrieval, where combining complementary modalities can reveal patterns and insights that remain hidden when each modality is independently processed [4]. Traditional methods typically processes each modality separately and then merges the results using simple concatenation or fixed-weight averaging techniques [2], but these basic fusion approaches often fail to capture the intricate interdependencies and correlations among modalities, resulting in representations that do not fully utilize the available information and consequently, limit the overall model performance [5].
@@
Figure 1 clearly illustrates the challenges inherent in multi-modal data: the first image (Figure 1(a)) demonstrates the complex interplay between textual cues and visual context, while the second image (Figure 1(b)) emphasizes the subtle visual nuances and layered textual information that must be interpreted accurately. These examples highlight the inherent difficulties faced by traditional fusion methods, which often rely on static, simplistic approaches that fail to capture the dynamic, context-dependent relationships between modalities. In response to this problem, researchers have turned to advanced pre-trained models to overcome these limitations. Models such as BERT have significantly advanced natural language processing [6], and convolutional neural networks have transformed computer vision [7, 8]. Although these models excel at extracting robust modality-specific features, integrating these powerful representations into a unified multi-modal framework remains challenging. Current fusion strategies often struggle to align and combine the complementary features effectively, underscoring the need for more adaptive, context-aware mechanisms such as those proposed in our Co-AttenDWG architecture to fully leverage the potential of multi-modal learning [9].
@@
To address these limitations, we propose a novel multi-modal Co-AttenDWG architecture that integrates dual-path encoding, co-attention with dimension-wise gating, and advanced expert fusion into a unified framework. Our model begins by projecting features from both modalities into a common embedding space, setting the stage for deep interaction. A dedicated co-attention mechanism then facilitates simultaneous, fine-grained cross-modal interactions. By incorporating a dimension-wise gating network, Co-AttenDWG dynamically modulates feature contributions at the channel level, ensuring that the most informative aspects of each modality are selectively emphasized during the fusion process. We validate our approach on challenging datasets such as MIMIC and SemEval Memotion 1.0, which demand robust multi-modal understanding. Experimental results reveal that our architecture significantly improves cross-modal alignment and outperforms state-of-the-art methods, demonstrating its potential to advance the field of multi-modal learning. This work not only addresses the current limitations of static fusion strategies but also opens avenues for more adaptive and context-aware integration of heterogeneous data sources in future research.
The key contributions of this work are summarized as follows:
  • We design a novel multi-modal Co-AttenDWG architecture that processes text and image features in a dual-path encoding framework, enabling robust alignment and refined representations.
@@
  • We introduce a co-attention mechanism with dimension-wise gating that dynamically modulates channel-level feature contributions, enhancing fine-grained cross-modal interactions.
@@
  • We develop an advanced expert fusion module that integrates learned gating with additional self-attention, effectively combining modality-specific representations into a unified embedding.
@@
The rest of this paper is organized as follows. Section II iscusses related work in multi-modal offensive content detection and cross-modal fusion techniques. Section III outlines the Co-AttenDWG framework, including its key components and architectural design. Section IV presents our experiment, results, and provides a detailed analysis. Finally, Section V concludes the paper and highlights future research directions.
II Literature
@@ III Proposed Methodology
Figure 2: A high-level overview of our Co-AttenDWG architecture for multi-modal offensive content detection. The image branch (top) processes memes through a pre-trained CNN (VGG16), extracting high-level visual features $I$. Meanwhile, the text branch (bottom) encodes input sentences using a language model such as DistilBERT, yielding textual features $T$. Both sets of features are projected into a shared embedding space and enter the co-attention gating block, where dimension-wise gating adaptively emphasizes salient channels from each modality. The attention outputs, highlighted as $A_{i}$ for images and $A_{t}$ for text, pass into a mixture-of-experts fusion mechanism that combines relevant cross-modal cues. Next, the fused representation flows through the Mamba block, which integrates local convolutional operations and multi-head self-attention to refine context. Finally, the aggregated features are projected via linear layers to produce the final output, representing the predicted offensive content class. This design promotes dynamic cross-modal interactions and expert gating, enabling effective offensive content detection in both text and images.
Figure 2 presents our proposed architecture, Co-AttenDWG, which addresses the challenges of multi-modal offensive content detection. Offensive content detection in multi-modal settings is challenging because text $X_{\text{text}}\in\mathbb{R}^{B\times L}$ and images $X_{\text{img}}\in\mathbb{R}^{B\times H\times W\times C}$ are processed through distinct pipelines that produce heterogeneous feature representations. For example, text is encoded using a model such as BERT [6] that generates hidden states $H\in\mathbb{R}^{B\times L\times D_{\text{text}}}$ and extracts the representative [CLS] token $h_{\text{CLS}}=H[:,0,:]\in\mathbb{R}^{B\times D_{\text{text}}},$ which is then projected into a common embedding space to obtain $T\in\mathbb{R}^{B\times D}.$ Similarly, image features are extracted from a CNN such as VGG16 [7] to produce a feature vector $f\in\mathbb{R}^{B\times D_{\text{img}}},$ which is also projected into the same space as $I\in\mathbb{R}^{B\times D}.$ Traditional fusion strategies, such as simple concatenation $F=[T;I]$, do not capture the dynamic, context-dependent cross-modal interactions needed for robust detection. To overcome these limitations, our method defines an adaptive fusion function $\mathcal{F}(T,I)$ that produces a unified feature representation $E\in\mathbb{R}^{B\times D}$ by aligning heterogeneous modalities. This figure illustrates how our approach, through bidirectional fusion with co-attentive dimension-wise gating and expert fusion, emphasizes the most relevant features from each modality to enhance detection performance.
@@ III-A Multi-Modal Feature Extraction and Projection
In the text branch, we tokenize the input and process it with a pre-trained model such as BERT [6] to obtain a sequence of hidden states, from which we extract the [CLS] token as an overall semantic summary. This token is then projected into a common $D$-dimensional embedding space using a learned linear transformation and reshaped into a sequence for later integration. In the image branch, each image is passed through a CNN like VGG16 [7] to extract high-level visual features, which are subsequently projected into the same $D$-dimensional space and reshaped into a sequence. Aligning both modalities in this shared embedding space reduces heterogeneity and lays a robust foundation for subsequent fusion processes that effectively combine complementary information for improved offensive content detection.
@@ III-D Expert Fusion
Dataset Description Original Balanced
MIMIC Non Misogynistic 2497 2497
Misogynistic 2409 2497
Memotion not_offensive 2657 2657
slight 2536 2657
very_offensive 1424 2657
hateful_offensive 213 2657
Note: The Memotion dataset mapping is {”not_offensive”: 0, ”slight”: 1, ”very_offensive”: 2, ”hateful_offensive”: 3}.
@@
TABLE II: Baseline Performance on MIMIC and Memotion Datasets (Accuracy and Macro F1 in %)
Model MIMIC (%) Memotion (%)
Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$)
mBERT [6] 77.98 77.59
BERT [6] 78.60 78.82
DistilBERT [32] 75.71 75.78
VGG16 [7] 84.58 84.58 76.83 76.78
ResNet50 [33] 84.68 84.56 77.36 77.25
EfficientNetV2 [34] 81.68 81.65 62.14 61.20
XLM-RoBERTa [35] 49.55 43.13
mBERT-VGG16 85.19 85.16
mBERT-ResNet50 85.99 85.98
mBERT-Efficient 83.08 82.90
RoBERTa + ResNet50 86.29 86.29
RoBERTa + VGG16 85.09 85.09
RoBERTa + Efficient 82.28 82.24
mCLIP [36] 85.79 85.78 82.60 82.66
VisualBERT [37] 86.39 86.39 81.28 81.26
BERT-ResNet50 82.08 82.00
BERT-Efficient 79.21 78.94
BERT-VGG16 81.10 81.00
DistilBERT-ResNet50 81.28 81.03
DistilBERT-VGG16 81.14 81.07
DistilBERT-Efficient 51.98 49.73
Co-AttenDWG 87.19 87.16 84.29 84.26
Improvements (+0.80$\uparrow$) (+0.77$\uparrow$) (+1.69$\uparrow$) (+1.60$\uparrow$)
Note: RoBERTa = XLM-RoBERTa [35], Efficient = EfficientNetV2 [34].

### v2

@@ Abstract
Multi-modal learning has emerged as a crucial research direction, as integrating textual and visual information can substantially enhance performance in tasks such as classification, retrieval, and scene understanding. Despite advances with large pre-trained models, existing approaches often suffer from insufficient cross-modal interactions and rigid fusion strategies, failing to fully harness the complementary strengths of different modalities. To address these limitations, we propose Co-AttenDWG, co-attention with dimension-wise gating, and expert fusion. Our approach first projects textual and visual features into a shared embedding space, where a dedicated co-attention mechanism enables simultaneous, fine-grained interactions between modalities. This is further strengthened by a dimension-wise gating network, which adaptively modulates feature contributions at the channel level to emphasize salient information. In parallel, dual-path encoders independently refine modality-specific representations, while an additional cross-attention layer aligns the modalities further. The resulting features are aggregated via an expert fusion module that integrates learned gating and self-attention, yielding a robust unified representation. Experimental results on the MIMIC and SemEval Memotion 1.0 datasets show that Co-AttenDWG achieves state-of-the-art performance and superior cross-modal alignment, highlighting its effectiveness for diverse multi-modal applications.
@@ I Introduction
Multi-modal learning has emerged as a transformative paradigm in artificial intelligence, driven by the necessity to integrate diverse data sources such as text, images, audio, and video to provide a holistic understanding of complex real-world scenarios [1, 2, 3]. This integration is particularly vital in tasks such as classification, sentiment analysis, and information retrieval, where the combination of complementary modalities reveals patterns and insights that remain hidden when each modality is processed independently [4]. Traditional methods typically process each modality separately and merge the results through simple concatenation or fixed-weight averaging [2]. However, such basic fusion approaches often fail to capture the intricate interdependencies and correlations among modalities, resulting in suboptimal representations and limiting overall model performance [5].
@@
Figure 1 exemplifies the multifaceted challenges inherent in multi-modal data integration. In Figure 1(a), the meme combines textual humor with a richly nuanced visual context, where accurate interpretation depends not only on the literal meaning of the text but also on subtle visual cues such as facial expressions, gestures, and background elements that add layers of meaning and sentiment. Similarly, Figure 1(b) presents a meme where layered textual cues interact with complex visual themes, including political symbolism and social context, requiring a sophisticated, fine-grained understanding of both modalities to fully grasp the intended message. These examples underscore the fundamental limitations of traditional static and simplistic fusion approaches that typically aggregate modalities without modeling their dynamic, context-dependent relationships. Such methods often fail to capture cross-modal dependencies and complementary information, resulting in suboptimal and sometimes misleading representations.
To address these challenges, recent research has leveraged powerful pre-trained models like BERT [6] and its multilingual and domain-adapted variants [7, 8] for deep language understanding, alongside convolutional neural networks [9, 10] and vision transformers [11] for visual feature extraction. While these architectures excel at generating robust, modality-specific embeddings, integrating them effectively remains a significant hurdle. Existing fusion strategies often rely on fixed or shallow combination mechanisms such as concatenation [2], early or late fusion [1], or simple attention mechanisms [12]. However, these approaches inadequately align and reconcile the heterogeneous representations from different modalities. Recent advances propose more sophisticated cross-modal attention and co-attention mechanisms that dynamically model inter-modal interactions at multiple granularities [13, 14, 15], alongside gating networks that adaptively weigh features to suppress noise and highlight complementary signals [16]. Transformer-based fusion modules [17] and graph neural networks for multimodal reasoning [18] have also shown promise in enhancing cross-modal alignment. These works highlight the growing consensus on the need for adaptive, context-aware fusion mechanisms capable of dynamically regulating cross-modal interactions at a fine-grained level. Such approaches selectively emphasize the most informative features from each modality depending on context, thereby improving interpretability and accuracy, particularly in complex tasks such as offensive content detection [19], sentiment analysis [20], and multi-modal reasoning [21].
@@
To address these shortcomings, we propose Co-AttenDWG, a novel multi-modal architecture that combines dual-path encoding with a co-attention mechanism enhanced by dimension-wise gating and advanced expert fusion. Our approach projects text and image features into a shared embedding space, where simultaneous, fine-grained cross-modal interactions occur via the co-attention mechanism. The dimension-wise gating network dynamically modulates channel-level feature contributions, selectively emphasizing the most informative components during fusion. We validate our approach on challenging datasets including MIMIC and SemEval Memotion 1.0, which require robust multi-modal comprehension. Experimental results demonstrate significant improvements in cross-modal alignment and state-of-the-art performance, illustrating the effectiveness and generalizability of our model.
The key contributions of this work are as follows:
  • We design a dual-path Co-AttenDWG architecture that robustly aligns and refines multi-modal representations.
@@
  • We introduce a dimension-wise gated co-attention mechanism to enable adaptive, fine-grained cross-modal interactions.
@@
  • We develop an expert fusion module that combines learned gating with self-attention to produce a unified, discriminative embedding.
@@
The rest of this paper is organized as follows. Section II iscusses related work in multi-modal offensive content detection and cross-modal fusion techniques. Section III outlines the Co-AttenDWG framework, including its key components and architectural design. Section IV presents our experiment, results, and provides a detailed analysis. Section V discusses the limitations of our study and prospective improvements that can be addressed in the future. Finally, Section VI concludes the paper and highlights future research directions.
II Literature Review
@@
Figure 2: A high-level overview of our Co-AttenDWG architecture for multi-modal offensive content detection. The image branch (top) processes memes through a pre-trained CNN (ResNet50), extracting high-level visual features $I$. Meanwhile, the text branch (bottom) encodes input sentences using a language model such as BERT, yielding textual features $T$. Both sets of features are projected into a shared embedding space and enter the co-attention gating block, where dimension-wise gating adaptively emphasizes salient channels from each modality. The attention outputs, highlighted as $A_{i}$ for images and $A_{t}$ for text, pass into a mixture-of-experts fusion mechanism that combines relevant cross-modal cues. Next, the fused representation flows through the Mamba block, which integrates local convolutional operations and multi-head self-attention to refine context. Finally, the aggregated features are projected via linear layers to produce the final output, representing the predicted offensive content class. This design promotes dynamic cross-modal interactions and expert gating, enabling effective offensive content detection in both text and images.
@@ III Proposed Methodology
Figure 2 presents our proposed architecture, Co-AttenDWG, which addresses the challenges of multi-modal offensive content detection. Offensive content detection in multi-modal settings is challenging because text $X_{\text{text}}\in\mathbb{R}^{B\times L}$ and images $X_{\text{img}}\in\mathbb{R}^{B\times H\times W\times C}$ are processed through distinct pipelines that produce heterogeneous feature representations. For example, text is encoded using a model such as BERT [6] that generates hidden states $H\in\mathbb{R}^{B\times L\times D_{\text{text}}}$ and extracts the representative [CLS] token $h_{\text{CLS}}=H[:,0,:]\in\mathbb{R}^{B\times D_{\text{text}}},$ which is then projected into a common embedding space to obtain $T\in\mathbb{R}^{B\times D}.$ Similarly, image features are extracted from a CNN such as ResNet50 [10] to produce a feature vector $f\in\mathbb{R}^{B\times D_{\text{img}}},$ which is also projected into the same space as $I\in\mathbb{R}^{B\times D}.$ Traditional fusion strategies, such as simple concatenation $F=[T;I]$, do not capture the dynamic, context-dependent cross-modal interactions needed for robust detection. To overcome these limitations, our method defines an adaptive fusion function $\mathcal{F}(T,I)$ that produces a unified feature representation $E\in\mathbb{R}^{B\times D}$ by aligning heterogeneous modalities. This figure illustrates how our approach, through bidirectional fusion with co-attentive dimension-wise gating and expert fusion, emphasizes the most relevant features from each modality to enhance detection performance.
@@ III-A Multi-Modal Feature Extraction and Projection
In the text branch, we tokenize the input and process it with a pre-trained Transformer such as BERT [6], which has been demonstrated to excel at capturing long-range dependencies and contextual semantics in language. We extract the [CLS] hidden state as a global summary token, then project it into a common $D$-dimensional embedding space via a learned linear layer and reshape it into a token sequence for downstream fusion. In the image branch, we employ a convolutional neural network (CNN) such as ResNet50 [10] to extract hierarchical visual features. CNNs continue to be the de facto standard for picture encoding because of their effective inductive biases, such as weight sharing and local receptive fields, which allow for consistent training on sparse data and quick inference. Similarly, these visual elements are molded into a pseudo-sequence and projected onto the common $D$-dimensional space.
@@ III-B Bidirectional Fusion with Co-Attentive Dimension-Wise Gating
Dataset Class Description Original Count Balanced Count
MIMIC Non-Misogynistic 2497 2497
Misogynistic 2409 2497
Non-Humiliation 4537 4537
Humiliation 369 4537
Non-Objectification 3462 3462
Objectification 1444 3462
Non-Prejudice 4032 4032
Prejudice 874 4032
Memotion not_offensive 2657 2657
slight 2536 2657
very_offensive 1424 2657
hateful_offensive 213 2657
Note: The Memotion (Offensive Content) dataset mapping is {”not_offensive”: 0, ”slight”: 1, ”very_offensive”: 2, ”hateful_offensive”: 3}.
TABLE II: Selected hyperparameters and data preparation details for Co-AttenDWG experiments.
Hyperparameter / Setting Option Best
Datasets MIMIC / Memotion / Both Both
Language Coverage English / Multilingual Hindi-English Both
Multilingual Models mBERT, XLM-RoBERTa mBERT, XLM-RoBERTa
Optimizer AdamW / SGD / RMSProp AdamW
[…]
Epochs 10 / 15 / 20 16 (Early Stop)
Learning Rate Scheduler None / Step / Dynamic Dynamic
Early Stopping 3 / 5 / 7 3
Attention Heads (Fusion) 4 / 8 / 12 8
Self-Attention Heads (Refinement) 2 / 4 / 8 4
MambaFormer Kernel Size 3 / 5 / 7 3
MambaFormer Depth 2 / 4 / 6 2
Dropout Rate 0.0 / 0.1 / 0.2 0.1
Text Tokenizer BERT / XLM-R / mBERT BERT / XLM-R
Image Normalization Standard / MinMax / None Standard (mean/std)
Image Size (MIMIC) $\times 160$ 160 / $\times 200$ 200 $\times 200$ 200 px
Image Size (Memotion) $\times 128$ 128 / $\times 160$ 160 $\times 160$ 160 px
@@ III-D Expert Fusion
TABLE III: Performance comparison of baselines and multimodal models on the MMIC and Memotion datasets. Best results in each column are bolded; second-best are underlined. Inference time is measured on a single NVIDIA RTX 2060 12GB GPU, batch size 1.
Model Time (ms) Misogyny (%) Objectification (%) Prejudice (%) Humiliation (%) Memotion (%)
Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
mBERT [6] 10.2 77.98 77.59 86.86 86.86 90.33 90.31 94.71 94.71
BERT [6] 10.2 78.60 78.82
DistilBERT [44] 6.1 75.71 75.78
VGG16 [45] 12.5 84.58 84.58 92.51 92.51 95.04 95.04 97.23 97.23 76.83 76.78
ResNet50 [10] 7.8 84.68 84.56 91.01 91.00 94.79 94.79 96.87 96.87 77.36 77.25
EffNetV2 [46] 6.5 81.68 81.65 90.90 90.90 94.23 94.23 96.99 96.99 62.14 61.20
XLM-R [47] 13.2 49.55 43.13 84.12 84.10 86.48 86.48 92.33 92.32
mBERT-VGG16 22.4 85.19 85.16 94.08 94.07 95.35 95.35 97.32 97.32
mBERT-ResNet50 17.5 85.99 85.98 94.51 94.51 96.09 96.09 97.28 97.28
mBERT-Efficient 16.3 83.08 82.90 93.65 93.65 95.29 95.29 97.56 97.56
RoBERTa+ResNet50 19.6 86.29 86.29 93.33 93.33 94.11 94.11 98.11 98.11
RoBERTa+VGG16 23.5 85.09 85.09 92.44 94.44 94.23 94.23 97.88 97.88
RoBERTa+EffNetV2 18.4 82.28 82.24 91.41 91.41 94.17 94.17 97.39 97.39
mCLIP [48] 16.0 85.79 85.78 94.73 94.73 95.06 95.05 98.90 98.90 82.60 82.66
VisualBERT [14] 17.2 86.39 86.39 94.51 94.51 97.02 97.02 98.91 98.91 81.28 81.26
ALBEF [49] 22.5 85.60 85.60 94.05 94.05 96.41 96.40 98.60 98.60 82.23 82.11
BLIP [50] 22.8 85.75 85.75 94.27 94.27 96.67 96.66 98.65 98.65 82.38 82.34
BERT-ResNet50 17.5 82.08 82.00
BERT-Efficient 16.3 79.21 78.94
BERT-VGG16 22.4 81.10 81.00
DistilBERT-ResNet50 13.9 81.28 81.03
DistilBERT-VGG16 18.6 81.14 81.07
DistilBERT-Efficient 13.2 51.98 49.73
Co-AttenDWG 31.1 87.19 87.16 94.80 94.80 97.15 97.15 98.80 98.80 84.29 84.26
Improvements +0.80$\uparrow$ +0.77$\uparrow$ +0.07$\uparrow$ +0.07$\uparrow$ +0.13$\uparrow$ +0.13$\uparrow$ -0.11$\downarrow$ -0.11$\downarrow$ +1.69$\uparrow$ +1.60$\uparrow$
Note: RoBERTa = XLM-RoBERTa [47], Efficient = EfficientNetV2 [46]. Inference time is measured on an NVIDIA RTX 2060 12GB GPU, batch size 1. Memotion (Offensive Content)
TABLE IV: Combinatorial ablation study of the Co-AttenDWG model on the MMIC and Memotion datasets. Each row disables or modifies one or more major components. Best results are bolded.
Model Variant Misogyny (%) Objectification (%) Prejudice (%) Humiliation (%) Memotion (%)
Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$) Acc ($\uparrow$) F1 ($\uparrow$)
w/o EF 83.58 83.56 93.86 93.86 88.03 87.98 97.80 97.80 80.90 80.84
w/o CA 85.59 85.57 93.50 93.50 91.82 91.81 98.10 98.10 81.19 81.01
w/o XA 83.58 83.57 94.01 94.01 92.75 92.74 97.60 97.60 79.12 79.26
w/o MF 84.98 84.95 93.29 93.29 90.95 90.93 97.85 97.85 81.19 81.31
w/o FF 85.27 85.25 93.78 93.78 91.40 91.38 97.90 97.90 80.81 80.69
w/o EF+MF 83.10 83.06 93.05 93.04 88.89 88.86 97.40 97.40 79.55 79.62
w/o CA+XA 82.21 82.20 93.13 93.12 88.41 88.39 97.10 97.10 78.66 78.59
w/o EF+CA+MF 81.40 81.32 92.57 92.55 87.07 87.03 96.80 96.80 77.12 77.04
2Heads 85.66 85.64 93.99 93.99 91.78 91.77 97.50 97.50 80.63 80.58
w/o MF+FF 82.71 82.68 93.00 93.00 89.18 89.17 97.55 97.55 79.89 79.87
Full (Co-AttenDWG) 87.19 87.16 94.80 94.80 97.15 97.15 98.80 98.80 84.29 84.26
EF = ExpertFusion; CA = Co-Attention; XA = Cross-Attention; MF = MambaFormer; FF = Fine-grained Fusion; 2Heads = Reduced Attention Heads (4$\rightarrow$2). Memotion (Offensive Content)
TABLE V: Impact of core architectural hyperparameters (number of experts, cross-attention heads, co-attention heads, MambaFormer kernel size, depth, dropout, pixel value, and learning rate) on macro F1 (%) for each label. Results are on the MIMIC and Memotion validation sets. Best per column are in bold.
# Experts Cross- Co- Kernel Depth Dropout Pixel LR Misogyny (%) Object. (%) Prejudice (%) Humil. (%) Memotion (%)
Attn Attn Size Value
4 4 4 7 4 0.10 224 $2\times 10^{-5}$ 85.11 93.00 95.01 97.21 80.98
8 4 4 9 4 0.15 200 $3\times 10^{-5}$ 85.69 93.22 95.33 97.48 81.31
8 8 4 5 6 0.20 160 $1\times 10^{-5}$ 86.13 93.60 95.68 97.93 81.98
8 8 8 7 8 0.05 128 $5\times 10^{-5}$ 86.79 94.10 96.92 98.51 84.01
8 8 8 11 6 0.10 200 $4\times 10^{-5}$ 86.52 94.00 96.30 98.34 83.45
8 8 8 5 3 0.30 128 $1.5\times 10^{-5}$ 86.35 94.22 96.20 98.22 83.01
8 8 4 3 2 0.10 200/160 $2\times 10^{-5}$ 87.16 94.80 97.15 98.80 84.26
8 8 8 7 4 0.00 224 $2.5\times 10^{-5}$ 86.88 94.35 96.75 98.08 83.99
Object. = Objectification, Humil. = Humiliation, Memotion (Offensive Content)
TABLE VI: Co-AttenDWG performance using different backbone models on the MIMIC and Memotion datasets (Accuracy and F1 in %). Best results are bolded. Memotion (Offensive Content)
Backbone Misogyny (%) Objectification (%) Prejudice (%) Humiliation (%) Memotion (%)
Text Image Acc F1 Acc F1 Acc F1 Acc F1 Acc F1
mBERT ResNet50 85.29 85.26 94.08 94.08 96.09 96.09 98.80 98.80
mBERT VGG16 85.29 86.83 94.73 94.73 95.23 95.23 98.80 98.80
mBERT EfficientNetV2 85.28 85.27 94.30 94.30 94.07 94.05 96.99 96.99
XLM-RoBERTa ResNet50 87.79 87.83 94.80 94.80 97.15 97.15 98.80 98.80
XLM-RoBERTa VGG16 86.17 86.82 91.91 91.91 97.02 97.02 98.80 98.80
XLM-RoBERTa EfficientNetV2 82.58 82.86 90.40 90.40 92.25 92.25 97.01 97.01
BERT ResNet50 84.29 84.26
BERT VGG16 83.66 83.61
BERT EfficientNetV2 81.19 81.01
DistilBERT ResNet50 81.81 81.81
DistilBERT VGG16 82.09 82.07
DistilBERT EfficientNetV2 79.97 80.01